If last year a brownie sale made $240.00, this year it made $360.00. what percent increase did it make this year.

360-240=20 20/100*360=72 Ans:72%

Answer by jim_thompson5910(35256)

First find the difference. Then divide that difference by the original amount.

360 - 240 = 120

120/240 = 0.5

0.5 ---> 0.5*100 = 50%

-------------------------------------------------------

Percent Increase: 50%